OpenStudy (anonymous):

Find the coordinates of a midpoint, AB.. A (12,-7) & B (-4,7)

OpenStudy (campbell_st):

midpoint... is the average of the x's and average of the y's x = (12 +(-4))/2 = 4 y(-7 + 7)/2 = 0

OpenStudy (anonymous):

k...

OpenStudy (anonymous):

@Kreshnik i tried .

OpenStudy (anonymous):

what have you tried ? :P

OpenStudy (anonymous):

to understand . See he didnt even show me how lol .

OpenStudy (anonymous):

yes he showed you how... he even did it for you ! ok the formula is \[\LARGE \left( {\frac{{{x_1} + {x_2}}}{2},\frac{{{y_1} + {y_2}}}{2}} \right)\] now you have: x_1=12 x_2=-4 and y_1=-7 y_2=7 substituting you'll have the above result